40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 290     Entry time: Thu Nov 30 12:18:41 2017
Author: Stephen 
Type: General 
Category: General 
Subject: Preparation for Modal Testing on 4 December 

Norna Robertson, Stephen Appert ||  29 Nov 2017, 2 pm to 4 pm  ||  227 Downs, CIT

We made some preparations for modal testing, but did not have enough time to make measurements. Below is an after-the-fact log, including some observations and photos of the current state of the OMC bench.

  1. Previous testing results at T1700471 (technical note in progress as of 30 Nov 2017).
    1. One goal of the next round: add damping material to equate with damping material of T1600494.
    2. Second goal of the next round: use a more localized sweep to better resolve the body mode around 1080 Hz -1100 Hz
  2. Transport Fixture was opened without issue, revealing the "Top" (suspending and cable routing) surface of the bench. Damping stacks were still in place from previous testing
  3. We removed the bolts from the damper stacks, but found that all masses with metal-viton interfaces had adhered to viton washers, causing the stacks to stick together.
    1. By using an allen key as a lever to wedge apart bottom mass and the bracket where they were joined by a viton washer, we separated the masses from the bracket.
    2. An allen key was used as a lever to push apart the two masses, which were also joined by a viton washer
    3. Once exposed, viton washers were pried from metal surfaces.
  4. After the damper stacks had been detached from the  No viton washer appeared to leave any residue or particulate - the separated parts all appeared as clean as they had been at the onset.
ELOG V3.1.3-